OpenStudy (liizzyliizz):

derivative of f(x)= cotx/sinx im kind of stuck using the quotient rule on this one...

OpenStudy (anonymous):

Do you mean cos x / sin x? If so, then you can just use substitution cos x = t, sin x = n, then derive n/t.

OpenStudy (liizzyliizz):

ohh. im still lost either way, can u expand on that?

OpenStudy (anonymous):

The quotient rule is for f'(x) = n/t = (n * t') - (t * n') / t² Now use what I've read above, and try to fill this formula.

OpenStudy (liizzyliizz):

that i know, and i did it, but my book says the answer is (-1-cos^2x)/sin^3x i want to know how it got to that. (i mean clearly im having trouble remembering basic trig identities..)

OpenStudy (anonymous):

is the numerator cot x or cos x? I don't know how to come to that answer either...

OpenStudy (liizzyliizz):

the numerator is cotx. ugh ive been sitting here for a while wondering how my textbook got that answer -_- It seems like it would be simple, and im just forgetting something

OpenStudy (anonymous):

start with the derivative of cot x? I don't know what it is, you could google it.

OpenStudy (ybarrap):

f(x)= cotx/sinx = cotx * csc x, since 1/sinx = cscx d/dx(cot(x) csc(x)) Use the product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = cot(x) and v = csc(x): = cot(x) (d/dx(csc(x)))+csc(x) (d/dx(cot(x))) The derivative of cot(x) is -csc^2(x): = cot(x) (d/dx(csc(x)))+csc(x) (-csc^2(x)) The derivative of csc(x) is -cot(x) csc(x): = cot(x) (-cot(x) csc(x))-csc^3(x) = -csc(x)(cot^2(x) + csc^2(x)) = -csc(x)(cos^2(x)/sin^2(x) + csc^2(x)) = -csc(x)(cos^2(x)csc^2(x) + csc^2(x)) = -csc^3(x)(cos^2(x) + 1) = (-1 - cos^2(x))csc^3(x) = (-1 - cos^2(x)/sin^3(x) Ans. (-1 - cos^2(x))/sin^3(x)

OpenStudy (jamesj):

Alternatively, f(x) = cot x / sin x = cos x /sin^2 x Hence applying the quotient rule: df/dx = (sin^2 x .(-sin x) - cos x. 2 sin x. cos x)/sin^4 x = -sin x ( sin^2 x + 2cos^2 x)/sin^4 x = -(1 + cos^2 x)/sin^3 x

OpenStudy (ybarrap):

Nice job James!